OpenStudy (anonymous):

On a trip to Italy, Cheyne rented a car to tour the country. As he turned onto the highway, he noticed that the speed limit was 95 km/h. How fast can Cheyne drive in mph?

OpenStudy (anonymous):

you need to convert km in miles 1 mile = 1.6 km ? miles = 95 km cross product and you get 95*1 = 1.6*? ? = 95/1.6 = 59.275 miles the unit of time is hour for both so you don't need any conversion for that so the answer is 59.275 mph

OpenStudy (anonymous):

thank you thank you thank you!

OpenStudy (anonymous):

your welcome! the hard part is know the conversion between km and miles, I believe it would be given in an exam but when you do your homework you can just google it if you are not sure.

OpenStudy (anonymous):

thats what i try to do but some of the links from google are hard to understand... thats how i found this website